OpenStudy (anonymous):

I need help ASAP, please!!!! I am stuck trying to do this integral integral of (-36x^(2)-69x-22)/(9x^(2)+18x+8) dx I been trying to do it using partial fractions but I am not getting anywhere T_T

rishavraj (rishavraj):

see divide the numerator by the denominator and express the integrand as \[Q(x) + \frac{ R(x) }{ 9x^2 + 18x + 8 }\]

OpenStudy (anonymous):

I am struggling solving for A & B :(

rishavraj (rishavraj):

hmmmmm why solve it using partial fractions.....

OpenStudy (anonymous):

I end up with 3x+10=A(3x+4)+B(3x+2)

rishavraj (rishavraj):

@lili18 see solve it this way.... divide the numerator with the denominator

OpenStudy (anonymous):

how do i solve for a and b on the following; 10=2a+4b? please, help!!!

rishavraj (rishavraj):

hmmmm u got 1 equation and 2 variables ???

OpenStudy (anonymous):

yes that when x=0

rishavraj (rishavraj):

to solve these type of question...... go on with the way I mentioned above

rishavraj (rishavraj):

see to solve it using partial fractions .... first step is divide the numerator by the denominator..... u will get \[- 4 + \frac{ 3x + 10 }{ 9x^2 + 18x + 8 }\] now solve using partial fractions :))

rishavraj (rishavraj):

so it would be \[\frac{ 3x + 10 }{ (3x + 2)(3x + 4) } = \frac{ A }{ 3x + 2 } + \frac{ B }{ 3x + 4 }\]

rishavraj (rishavraj):

@lili18 can u solve it now???

rishavraj (rishavraj):

@lili18 u there??

OpenStudy (dumbcow):

\[3A +3B = 3 \rightarrow A+B = 1\] \[4A +2B = 10 \rightarrow 2A +B = 5\] solve by substitution \[2A + (1-A) = 5\] \[A = 4\] \[B = -3\]